OpenStudy (anonymous):

Show that (√6 + √2)/(√6 - √2) = 2+√3

OpenStudy (anonymous):

(√6 + √2)/(√6 - √2) = 2+√3 ( x+ 1 ) ( x - 1 ) difference of squares (√6 + √2)/(√6 - √2) = 2+√3 | | _______________________ \do you see that this follows difference of squares format?

OpenStudy (anonymous):

yes I do see it as that .. but remember itz divided by right ..

OpenStudy (anonymous):

\[\frac{\sqrt{6}+\sqrt{2}}{\sqrt{6}-\sqrt{2}}\times \frac{\sqrt{6}+\sqrt{2}}{\sqrt{6}+\sqrt{2}}\] is probably a good start

OpenStudy (anonymous):

yes it is because it is a surd .. but when i work it out i end up with two root threes and itz suppose to be one so im a bit confused .

OpenStudy (anonymous):

same thing what satellite did.

OpenStudy (anonymous):